Grade upto college level Physical Chemistry

The standard molar heats of formation of ethane, carbon dioxide and liquid water are -21.1, -94.1 and -68.3 kcal respectively. Calculate the standard molar heat of combustion of ethane.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
The required chemical reaction.
2C2H6 + 7O2 → 4CO2 + 6H2O; ∆H = x
Note that since 2 moles of ethane are reacting, the ∆H of the reaction will be ½ x.
The thermochemical equations for the given data are written as below.
(i) C(s) + O2(g) → Co2(g); ∆H = - 94.1 kcal
(ii) H2(g) ½O2(g) → H2O(g); ∆H = - 68.3 kcal
(iii) 2C(s) + 3H2(g) → C2H6(g); ∆H = - 21.1 kcal
We know that ∆H = HProducts – HReactants
∆H = 4\Delta H_{CO_{2}} + 6\Delta H_{H_{2}O} - (2\Delta H_{C_{2}H_{6}} +7\Delta H_{O_{2}})
∆H = 4 * - 94.1 + 6 * - 68.3 – (2 * - 21.1 + 0)
= - 376.4 – 409.8 + 42.2 = - 744.0 kcal/2 mole of ethane
= -372.0 kcal/mole of ethane

Thanks
jitender
askIITians Faculty